# Online Decision-Focused Learning

### v1

@@ Appendix B An online decision-focused algorithm with static regret guarantees
2:  for each $t\in\{1,\ldots,T\}$ do
5:     Draw $\sigma_{t}\in\mathbb{R}^{d}$ such that for all $j\in[d]$, $\sigma_{j,t}\sim\mathrm{Exp}(\eta)$.
6:     Update
$\theta_{t+1}=\mathcal{O}_{\xi,\chi}\mathopen{}\mathclose{{}\left(\sum_{i=1}^{t}\tilde{f}_{i}-\mathopen{}\mathclose{{}\left\langle\sigma_{t},\cdot}\right\rangle}\right)$
7:  end for
In Algorithm 3, recall that $\tilde{f}_{i}$ are the surrogate losses defined in Equation 5. At each time $t\in[T]$, Algorithm 3 first makes a decision $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ based on the current parameters $\theta_{t}\in\Theta$. Then, it observes $\bar{g}_{t}(X_{t})\in\mathcal{Z}$, determines the surrogate objective $\tilde{f}_{t}$. Next, it updates $\theta_{t+1}$ following the FTPL algorithm applied to the surrogate losses $(\tilde{f}_{t})_{t\in[T]}$.
In order to obtain static regret bounds, we follow a similar route than in Section 4.1: we work on the simplex $\mathcal{W}_{0}$ and use the negative entropy $\mathcal{R}_{0}$ as regularizer to define $(\tilde{f}_{t})_{t\in[T]}$. We refer to Section 4.1 for a precise definition of these notions. We then have the following result.
Proposition 4.
Assume H 1, H 2 and having access to an ($\xi,\chi$ )-approximate optimization oracle $\mathcal{O}$. Fix for any $t$, $\alpha_{t}=\alpha$ in (6), $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and that DF-FTPL (Algorithm 3) is instantiated with $\eta>0$. Then, for any $\theta\in\Theta$:
$T^{-1}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{S}}(\theta)}\right]$ $\leqslant\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\eta m^{2}D\frac{1}{\alpha^{2}}+\frac{m(\chi T+D)}{\eta T}+\xi+\chi m\frac{1}{\alpha}+\alpha\ln(d)}\right)\;,$
Thus, it is possible to have a decision-focused algorithm with static regret guarantees. A notable strength of Algorithm 3 compared to Algorithm 1 is that it does not require any optimism to run. However, this comes at a major cost: a dependency in the dimension $m^{2}$ of $\Theta$ which is not involved in Proposition 1 thanks to the optimistic approach.
The proof of Proposition 4 is deferred to Section F.3.
Appendix C Additional material on our numerical experiment.
In this section, we provide more detail about the experimental setup in Section 5 and perform additional numerical simulations.
Details on the experimental setup.
We start by explaining more precisely how the data used in our experiment are drawn. First, for any $t\in[T]$, $X_{t}\in\mathbb{R}^{K\times p}$ is constructed as follows: (i) we generate a Toeplitz covariance matrix $\Sigma=(\rho^{\mid i-j\mid})_{(i,j)\in[K]^{2}}$ for some $\rho\in(0,1)$, (ii) apply a Cholesky decomposition to it: $\Sigma=LL^{\mathsf{\scriptscriptstyle T}}$, and (iii) generate a matrix $\bar{X}_{t}$ with standard Gaussian entries. Then, $X_{t}$ is defined as $X_{t}=L\bar{X}_{t}$. This introduces correlation between features, which makes convergence harder for an online linear model. Second, we generate $v_{t}\in\mathbb{R}^{K}$ as follows:
$v_{t}=\min\mathopen{}\mathclose{{}\left(\max\mathopen{}\mathclose{{}\left(\tilde{v}_{t},0}\right),1}\right)\quad\text{with}\quad\tilde{v}_{t}=\alpha\sin^{4}((2X_{t}\theta^{\star}_{t})^{-1})+\varepsilon\;.$
In the above equation, $\alpha>0$ is a constant, $\theta_{t}^{\star}$ follows the following autoregressive recursion:
$\theta^{\star}_{t}=\frac{1}{2}\theta^{\star}_{t-1}+\frac{1}{2}\zeta_{t}\,,\quad\text{where}\quad\zeta_{t}\sim\mathrm{N}(0,I_{K})\;,$
We now present the values used for the different parameters in our experiment. We consider $K=5$ items and $p=10$ features. The horizon is set to $T=2000$. For each plot, we run $N=70$ times DF-OGD and PF-OGD. The reported error bars are 95% Gaussian confidence intervals (from the CLT). In (12), $\alpha=45$ and the correlation coefficient for $X_{t}$ is $\rho=0.8$. Our algorithms are instantiated with the following parameters (optimized through a grid search procedure). First, DF-OGD runs with $(\alpha_{t})$ and $(\eta_{t})$ as in Corollary 1. The oracle from H 2 is obtained through a SGD algorithm performing $t_{\vartheta}=10$ steps at each iteration with a learning rate set to $\eta_{\vartheta}=0.01$, while being initialized at $\vartheta_{t-1}$. On the other hand, PF-OGD runs with a learning rate $\eta=0.01$. Our code mostly builds on pytorch and numpy.
Deviation from realizability.
As argued in Section 1, prediction-focused learning would be optimal if the prediction model made no mistake, since in this case problem (2) would perfectly align with (1). As a consequence, one would expect prediction-focused learning to perform very well in the realizable setting, that is when for any $t\in[T]$, there exists $\bar{\theta}_{t}\in\Theta$ such that $\bar{g}_{t}(X_{t})=g(\bar{\theta}_{t},X_{t})$. On the flip side, it is likely to struggle in a mispecified setting. This is in contrast with DF-OGD, whose theoretical performance analysis does not highlight any peculiar dependence on realizability. We therefore hypothetize that PFL should outperform DFL in the realizable case, and DFL should take the upper hand when the prediction model is highly mispecified. To test this hypothesis, we run an alternative experiment where we interpolate between the well-specified, linear case, and ill-specified, non-linear one. More precisely, for $\gamma\in[0,1]$, we generate $(X_{t},v^{(\gamma)}_{t})_{t\in[T]}$ as follows:
In Figure 3, we make $\gamma$ vary from 0 to 1. At each value, we plot the average cumulated cost gap between DF-OGD and PF-OGD, averaged over $50$ runs with horizon $T=1000$. This experiment provides empirical support to the previous conjecture, namely, DFL becomes a competitive option when the prediction model is ill-specified.
Figure 3: Average cumulated cost gap PFL - DFL as a function of $\gamma$.
Appendix D A stronger version of H 2 for a stronger notion of regret
In this section, we present a stronger version of H 2, displayed in H 4. We show that under this new assumption, our guarantees still hold for a stronger notion of regret.
Here, we focus on the simplex setting of Section 4.1: we take $\mathcal{W}=\mathcal{W}_{0}$ the simplex and $\mathcal{R}=\mathcal{R}_{0}$ the negative entropy. This section show that at the cost of slightly strengthening this assumption, we may obtain regret bounds for DF-OGD with a stronger version of regret, which we define below.
A novel dynamic regret.
We define the pseudo dynamic regret as, for any $t$ any loss sequence $(f_{t})_{t\in[T]}$,
$\mathfrak{R}_{T}^{\mathrm{d}}:=\sum_{t\in[T]}f_{t}(\theta_{t})-\sum_{t\in[T]}\inf_{\theta\in\Theta}f_{t}(\theta).$
This notion is exactly what appeared in [77]. Note that since $\mathbb{E}\mathopen{}\mathclose{{}\left[\inf_{\theta\in\Theta}h(\theta,X)}\right]\leqslant\inf_{\theta\in\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[h(\theta,X)}\right]$, we immediately obtain that:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right].$
In this sense, obtaining a guarantee on $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right]$ is stronger than on $\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$.
Strengthening H 2.
We propose the following assumption:
H 4.
There exist $C_{0}>0$ and $\beta\in\mathopen{}\mathclose{{}\left[0,1}\right]$, such that for any $t\in[T]$ and $\varepsilon\in\mathopen{}\mathclose{{}\left[0,1}\right]$,
$\mathbb{P}\mathopen{}\mathclose{{}\left(\forall\theta\in\Theta,\>\inf_{j\neq I_{t}(\theta)\;}u_{j}(\theta,X_{t})-u_{I_{t}(\theta)}(\theta,X_{t})\geqslant\varepsilon\,\mid\,\mathcal{H}_{t-1}}\right)\geqslant 1-C_{0}\varepsilon^{\beta}\;,$
This reinforcement of H 2 comes from the fact that we took $\forall\theta\in\Theta$ inside the probability and not outside. Based on this assumption we adapt Lemma 3 to get the following lemma.
Lemma 1.
Assume H 4, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $\theta\in\Theta$ and $t\in[T]$,
$\mathbb{E}\mathopen{}\mathclose{{}\left[\sup_{\theta\in\Theta}\mathopen{}\mathclose{{}\left\|\tilde{w}_{t}(\theta)-w_{t}^{\star}(\theta)}\right\|_{1}\mathopen{}\mathclose{{}\left|\mathcal{H}_{t-1}}\right.}\right]\leqslant\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)\;.$
Lemma 1 shows that assuming H 4 instead of H 2 transform the control of the supremum of an expectation of a supremum (Lemma 3) into one on the expectation of a supremum.
Exploiting Lemma 1, we prove a novel regret bound involving the deterministic dynamic regret.
Proposition 5.
Assume H 1, H 4, H 3, $\mathcal{W}=\mathcal{W}_{0}$, $\mathcal{R}=\mathcal{R}_{0}$ and that Algorithm 1 is instantiated with a non-increasing sequence $\mathopen{}\mathclose{{}\left(\eta_{t}}\right)_{t\in[T]}$. Then, denoting $P_{T}=\sum_{t=1}^{T-1}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|$,
$\frac{1}{T}\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}^{\mathrm{d}}}\right]$ $\leqslant\frac{D_{\Theta}^{2}+D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2T\eta_{T}}+\sum_{t\in[T]}\frac{25D_{\mathcal{Z}}G\eta_{t}}{32T\alpha_{t}^{2}}$
where $\mathbb{E}$ denotes the expectation on both data and the intrinsic randomness of DF-OGD.
From Proposition 5 we can recover the same convergence rate than Corollary 1 and also perform a similar study than in Section 4.2 to have results for a general polytope. The tradeoff highlighted is that strengthening H 2 into H 4 is the price to pay to control the stronger notion of regret to $\mathfrak{R}_{T}^{\mathrm{d}}$.
The proofs of Lemma 1 and Proposition 5 are deferred to appendix Section F.4.
Appendix E Regret bound with an approximate optimal solution.
@@ Appendix D A stronger version of H 2 for a stronger notion of regret
Assume the assumptions from Proposition 2 and that at each step, $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ is replaced by $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ which satisfies
@@
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]$ $\leqslant\frac{D_{\Theta}^{2}-D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{}\left[P_{T}}\right]}{2\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{}\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{}\left\|b-Aw}\right\|_{\infty}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}\right)^{2}$
@@
The proof Proposition 6 is deferred to Section F.5.
Appendix F Proofs.
F.1 Proofs of Section 4.1.
@@
Assume H 1, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}$-Lipchitz almost surely, with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$.
@@
Assume H 2, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $\theta\in\Theta$ and $t\in[T]$, for any $\theta\in\theta$
@@
In this proof, $I_{t}(\theta)$ from H 2 is denoted $I_{t}$ to lighten notation. We have by assumption:
@@
Moreover, for any $\varepsilon>0$ we have by H 2 and Lemma 4 that:
@@
Assume that for any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}-$Lipschitz almost-surely and that the sequence of steps $(\eta_{t})_{t\geqslant 1}$ is non-increasing. Denote $\tilde{\mathfrak{R}_{T}}=\sum_{t\in[T]}\tilde{f}_{t}(\theta_{t})-\tilde{f}_{t}(\vartheta_{t})$. Then it holds almost surely that:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\tilde{\mathfrak{R}}_{T}}\right]\leqslant\frac{1}{2\eta_{T}}\mathopen{}\mathclose{{}\left(D_{\Theta}^{2}+2D_{\Theta}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|}\right)+\sum_{t=1}^{T}\frac{\eta_{t}}{2}K_{t}^{2}\;,$
where the expectation is taken over the sequence $(u_{1},\ldots,u_{T})$ in Algorithm 1.
Proof.
[…]
where ~$\tilde{\nabla}(u_{t})=\nabla\tilde{f}_{t}(\vartheta_{t}+u_{t}(\theta_{t}-\vartheta_{t})$ as in Algorithm 1. To control (26) we note that by definition of $\theta_{t+1}$ in Algorithm 1,
where we have used that ~$\mathopen{}\mathclose{{}\left\|\tilde{\nabla}_{t}(u_{t})}\right\|^{2}\leqslant K_{t}^{2}$ because $\tilde{f}_{t}$ is $K_{t}-$Lipschitz. Re-arranging (27) and taking the expectation over $(u_{1},\ldots,u_{T})$ yields:
Now notice that, for any $t\in\{1,\ldots,T-1\}$,
$=\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}$
where we have used H 1-(i) and the reversed triangular inequality in the last line. It follows that:
$-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t}}\right\|^{2}$ $\leqslant-\mathopen{}\mathclose{{}\left\|\theta_{t+1}-\vartheta_{t+1}}\right\|^{2}+2D_{\Theta}\mathopen{}\mathclose{{}\left\|\vartheta_{t+1}-\vartheta_{t}}\right\|$ (30)
Plugging (30) into (F.1) then gives:
Thus, summing over $T$ and plugging the resulting sum in (26) gives:
where we used that for all $t$, $\mathopen{}\mathclose{{}\left\|\theta_{t}-\vartheta_{t}}\right\|\leqslant D_{\Theta}$ by H 1-(i) and that for all $t\in[T],\,\eta_{t-1}\geqslant\eta_{t}\geqslant\eta_{T}$. Then, we have by telescoping the first sum:
See 1
Proof.
Observe that the dynamic regret can be decomposed as follows:
$\mathfrak{R}_{T}$ $=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta_{t})-\tilde{w}_{t}(\theta_{t})}\right\rangle}\right]+\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle}\right]$ (31)
Taking the $\sup$ over $\Theta$, defining by $\mathrm{Reg}_{T}:=\sum_{t\in[T]}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),\tilde{w}_{t}(\theta_{t})-\tilde{w}_{t}(\vartheta_{t})}\right\rangle}\right]$ in (31), using H 3 in (32), and noticing that $\inf_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)]-\inf_{\theta\in\Theta}\mathbb{E}_{t}[\tilde{f}_{t}(\theta)]\leqslant\sup_{\theta\in\Theta}\mathbb{E}_{t}[f_{t}(\theta)-\tilde{f}_{t}(\theta)]$ in (33) leads to:
$\mathfrak{R}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\langle\bar{g}_{t}(X_{t}),w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\rangle}\right]+\xi T$
Using the fact that $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{\infty}$ is dual to $\mathopen{}\mathclose{{}\left\|\,\bm{\cdot}\,}\right\|_{1}$:
$\mathfrak{R}_{T}$ $\leqslant\mathrm{Reg}_{T}+2\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]+\xi T$
Since for any $t\in[T]$, $\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{\infty}\leqslant\mathopen{}\mathclose{{}\left\|\bar{g}_{t}(X_{t})}\right\|_{2}\leqslant D_{\mathcal{Z}}$ almost surely by H 1-(ii),
$\mathfrak{R}_{T}$ $\leqslant\mathrm{Reg}_{T}+2D_{\mathcal{Z}}\sum_{t\in[T]}\sup_{\theta\in\Theta}\mathbb{E}_{t}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\|w^{\star}_{t}(\theta)-\tilde{w}_{t}(\theta)}\right\|_{1}}\right]+\xi T\;$ (34)
First, notice that, by Lemma 3, for all $t$:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]\leqslant\mathbb{E}\mathopen{}\mathclose{{}\left[\mathrm{Reg}_{T}}\right]\\ +2D_{\mathcal{Z}}\sum_{t\in[T]}\alpha_{t}\mathopen{}\mathclose{{}\left(1+2\ln(d)C_{0}\mathopen{}\mathclose{{}\left(\ln\mathopen{}\mathclose{{}\left(\frac{2}{\alpha_{t}}}\right)+(1-\beta)\ln^{2}(\alpha_{t}\ln(d))}\right)}\right)+\xi T\;.$ (36)
To bound the first term in the right-hand side of Equation 36. Remark that, by the definition of conditional expectation:
One recognises the definition of $\tilde{\mathfrak{R}_{T}}$ of Lemma 5. Thus, by this lemma, we know that for any $t\in[T]$, because $\tilde{f}_{t}$ is $K_{t}-$Lipschitz with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$ almost surely by Lemma 2, we have:
Plugging Equation 37 into Equation 36 and then dividing by $T>0$ on both sides gives the desired result.
See 1
Proof.
We know by Proposition 1 that for any decreasing sequence of learning rates $(\eta_{1},\ldots,\eta_{T})\in\mathbb{R}_{+}^{T}$ and any regularization coefficients $(\alpha_{1},\ldots,\alpha_{T})\in\mathbb{R}_{+}^{T}$, the average regret of Algorithm 1 is bounded as follows:
$\mathbb{E}\mathopen{}\mathclose{{}\left[\mathfrak{R}_{T}}\right]=\tilde{\mathcal{O}}\mathopen{}\mathclose{{}\left(\mathbb{E}\mathopen{}\mathclose{{}\left[\frac{1+P_{T}}{\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{\alpha_{t}^{2}}+(1+\ln d)\sum_{t\in[T]}\alpha_{t}+\xi T}\right]}\right)\;.$
With $\alpha_{t}\propto t^{-1/4}(1+P_{t})^{1/4}$ and $\eta_{t}\propto\alpha_{t}t^{-1/2}(1+P_{t})^{1/2}\propto t^{-3/4}(1+P_{t})^{3/4}$, the sequence $(\eta_{t})_{t\in[T]}$ is non-increasing because so is $(t^{-1}(1+P_{t}))_{t\in[T]}$ by assumption. Moreover, the first term satisfies:
$\frac{1+P_{T}}{\eta_{T}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\frac{T^{3/4}}{(1+P_{T})^{3/4}}\mathopen{}\mathclose{{}\left(1+P_{T}}\right)}\right)\quad\text{so}\quad\frac{1+P_{T}}{T\eta_{T}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(\frac{1+P_{T}}{T}}\right)^{1/4}}\right)\;.$
For the second term,
$\sum_{t=1}^{T}\frac{\eta_{t}}{\alpha_{t}^{2}}=\sum_{t=1}^{T}\frac{\sqrt{1+P_{t}}}{\sqrt{t}\alpha_{t}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left(\frac{1+P_{t}}{t}}\right)^{1/4}}\right)\;..$
Moreover, since $1+P_{t}\leqslant 1+P_{T}$ for any $t\in[T]$, we have:
$\sum_{t=1}^{T}\frac{\eta_{t}}{\alpha_{t}^{2}}=\mathcal{O}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(1+P_{T}}\right)^{1/4}\sum_{t=1}^{T}\mathopen{}\mathclose{{}\left(\frac{1}{t}}\right)^{1/4}}\right)=\mathcal{O}\mathopen{}\mathclose{{}\left(\mathopen{}\mathclose{{}\left(1+P_{T}}\right)^{1/4}T^{3/4}}\right)\;,$
$\sum_{t\in[T]}\alpha_{t}\leqslant(1+P_{T})^{1/4}\sum_{t\in[T]}t^{-1/4}=\mathcal{O}((1+P_{T})^{1/4}T^{3/4})\;,$
and dividing by $T>0$ yields the desired rate. ∎
F.2 Proofs of Section 4.2.
Lemma 6.
@@
Assume H 1, $\mathcal{W}=\mathcal{W}_{1}$ and $\mathcal{R}=\mathcal{R}_{1}$. Then for any $t\in[T]$,$\tilde{f}_{t}$ is differentiable and for any $\theta\in\Theta$, $\nabla\tilde{f}_{t}(\theta)=\nabla\tilde{w}_{t}(\theta)\bar{g}_{t}(X_{t})$ where:
@@
Assume H 1 and $\mathcal{R}=\mathcal{R}_{1}$. For any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}$-Lipschitz almost-surely, with
@@
Assume H 2, $\mathcal{W}=\mathcal{W}_{1}$ and $\mathcal{R}=\mathcal{R}_{1}$. Then for any $t\in[T]$ and $\theta\in\Theta$,
@@
which is uniquely determined by strong convexity of the objective. We know by [15, page 566] that:

### v2

@@ Appendix E Additional experimental material
2:for each $t\in\{1,\dots,T\}$ do
3:  Observe $X_{t}$, predict $\hat{c}_{t}=g(\theta_{t},X_{t})$.
4:  Play $v^{\star}_{t}(\theta_{t})=\operatorname*{\arg\!\min}_{v\in\mathcal{\mathcal{V}}}\hat{c}_{t}^{\mathsf{\scriptscriptstyle T}}v$, observe $c_{t}$.
5:  Compute $v^{\star}(\hat{c}_{t})=\operatorname*{\arg\!\min}_{v\in\mathcal{\mathcal{V}}}\hat{c}_{t}^{\mathsf{\scriptscriptstyle T}}v$ and $v^{\star}(c_{t})=\operatorname*{\arg\!\min}_{v\in\mathcal{\mathcal{V}}}c_{t}^{\mathsf{\scriptscriptstyle T}}v$.
6:  Define SPO+ surrogate:
$\ell^{\mathrm{SPO+}}(\hat{c}_{t},c_{t})=2c_{t}^{\mathsf{\scriptscriptstyle T}}v^{\star}(\hat{c}_{t})-c_{t}^{\mathsf{\scriptscriptstyle T}}v^{\star}(c_{t})-\hat{c}_{t}^{\mathsf{\scriptscriptstyle T}}v^{\star}(\hat{c}_{t}).$
7:  Update $\theta_{t+1}=\Pi_{\Theta}\big(\theta_{t}-\eta\,\nabla_{\theta}\ell^{\mathrm{SPO+}}(\hat{c}_{t},c_{t})\big)$, where the gradient is computed via chain rule through $\hat{c}_{t}=g(\theta_{t},X_{t})$.
8:end for
Appendix F Regret bound with an approximate optimal solution.
@@
Assume the assumptions from Theorem˜2 and that at each step, $w^{\star}_{t}(\theta_{t})\in\mathcal{W}$ is replaced by $\underline{w}_{t}(\theta_{t})\in\mathcal{W}$ which satisfies
@@
$\mathbb{E}\mathopen{}\mathclose{{\left[\mathfrak{R}^{d}_{T}}}\right]$ $\leqslant\frac{D_{\Theta}^{2}-D_{\Theta}\mathbb{E}\mathopen{}\mathclose{{\left[P_{T}}}\right]}{2\eta_{T}}+\sum_{t\in[T]}\frac{\eta_{t}}{2}\mathopen{}\mathclose{{\left(\frac{GD_{\mathcal{Z}}\sup_{w\in\mathcal{W}_{1}}\mathopen{}\mathclose{{\left\|b-Aw}}\right\|_{\infty}}{\alpha_{t}\lambda_{\min}(AA^{\mathsf{\scriptscriptstyle T}})}}}\right)^{2}$
@@
The proof Proposition˜5 is deferred to Section˜G.6.
Appendix G Proofs
G.1 Proofs of preliminary lemmas.
G.1.1 Simplex
@@
Assume H˜1, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}$-Lipchitz almost surely, with $K_{t}=5D_{\mathcal{Z}}G(4\alpha_{t})^{-1}$.
@@
Assume H˜2, $\mathcal{W}=\mathcal{W}_{0}$ and $\mathcal{R}=\mathcal{R}_{0}$. Then for any $\theta\in\Theta$ and $t\in[T]$, for any $\theta\in\theta$
@@
In this proof, $I_{t}(\theta)$ from H˜2 is denoted $I_{t}$ to lighten notation. We have by assumption:
@@
Moreover, for any $\varepsilon>0$ we have by H˜2 and Lemma˜3 that:
@@
G.1.2 General polytope
In this section, we denote by $\mathcal{W}_{1}$ the general polytope described in LABEL:eq:polytope and $\mathcal{R}_{1}$ the log-barrier regularization described in Equation˜7.
@@ G.1.2 General polytope
Assume H˜1, $\mathcal{W}=\mathcal{W}_{1}$ and $\mathcal{R}=\mathcal{R}_{1}$. Then for any $t\in[T]$$\tilde{f}_{t}$ is differentiable and for any $\theta\in\Theta$, $\nabla\tilde{f}_{t}(\theta)=\nabla\tilde{w}_{t}(\theta)\bar{g}_{t}(X_{t})$ where:
@@
Assume H˜1 and $\mathcal{R}=\mathcal{R}_{1}$. For any $t\in[T]$, $\tilde{f}_{t}$ is $K_{t}$-Lipschitz almost-surely, with
@@
Assume H˜2, $\mathcal{W}=\mathcal{W}_{1}$ and $\mathcal{R}=\mathcal{R}_{1}$. Then for any $t\in[T]$ and $\theta\in\Theta$,
@@
which is uniquely determined by strong convexity of the objective. We know by [14, page 566] that: